# Examining Input Modalities and Visual Feedback Designs in Mobile Expressive Writing

### v1

@@ 5.1.2. Dependent Variables for Stress-relieving Effect
Similarly to our first study, we collected immediate stress change and post-experiment cathartic effect. But we decided not to collect the data related to the content of the expression as our focus was the feedback after the expression. We also obtained the following quantitative data for our analysis.
@@
To gauge how the participants’ stress levels changed over a longer period, we used Perceived Stress Scale (PSS) (Cohen et al., 1994). This assessment tool consists of 14 questions about how often each certain situation occurred in the past week on a five-point Likert scale ranging from (0: never – 4: very often). After collecting the score every week, we subtracted the after score from the before score and used it as the long-term stress change for the condition they experienced that week.
@@ 5.1.3. Dependent Variables for Usability and Preference
We collected ease of use, mental workload and preference in this study as well. In addition, we collected the following effort measurement. All four were collected from the post-experiment questionnaire.
Effort
To capture perceptions of the effort required, we asked “This feedback requires effort,” based on previous work compared the usability of digital coping methods (Howe et al., 2022). The response was in a five-point Likert scale (1: “Not at all” – 5:“Strongly agree”).
@@ 5.2. Procedure
The process in the onboarding and post-experimental sessions was the same as in Study 1. In this study, we explained our participants the objective of our study to examine the effects of three feedback designs at this time. The pre-experiment questionnaire included the questions from PSS for long-term stress change. The onboarding required approximately 30 minutes to complete on the first day. We split participants into six groups with considering the balance of gender, age, and occupations to fully counter-balance the order of the presentations of the three feedback conditions.
Participants are under each of the visual feedback conditions for each week of the experiment (Day 2–8, Day 9–15, and Day 16–22). On the mornings of days 9 and 16, we asked participants to fill out PSS to track their long-term stress level trends.
We conducted a post-experiment questionnaire once on the final day. We then asked the participants to fill out a post-experiment questionnaire (Table 4 in Appendix B), which included questions for PSS and about post-experiment cathartic effect, effort, ease of use, and mental workload of each feedback condition and the preference on feedback conditions. We also collected qualitative data with open-ended questions, asking participants for their thoughts or opinions about the positive and negative aspects of each feedback condition and their reasons for their preferences. We also asked when they would like to use each condition and what other feedback designs they would like to have. The post-experiment questionnaires took around 30 minutes to complete on the final day. We offered all participants approximately 21 USD in the local currency for the completion of the study.
We performed data analysis in the same way as we did in our first study.
@@ 5.3. Participants
We used the same crowdsourcing service as Study 1. We also followed the same procedure to choose participants. As a result, we recruited 36 participants (18 women, 18 men; PB1 – 36) in their 20s to 60s (Table 2). All participants self-claimed that they did not know of the existence of our first study.
@@ 6. Study 2 Results
We collected 550 expressive writing data with immediate stress change data, and the response rate was 99.4%. Among 550 expressive writing data, 186, 183, and 181 were under the None, Passive, and Interactive conditions, respectively.
@@ 6.1. Quantitative Results for Stress-relieving Effect
The average difference between pre- and post-experimental PSS scores were 0.39 ( = 5.23), -0.58 ( = 5.03), and -2.22 ( = 7.06) in the None, Passive, Interactive conditions, respectively (Figure 6). Our Friedman test did not reveal any significant difference ( = 1.22, = .54, = .117). For the Interactive condition, our Wilcoxon signed-rank test found a significant difference between pre- and post-experimental phases ( = 1, = 1.05, ¡ .05, = .175). As for the None and Passive conditions, our Wilcoxon signed-rank test did not reveal a significant difference between before and after (None: = 1, = 0.267, = .49, = 0.0445; Passive: = 1, = 0.0393, = .25, = 0.00655).
The average PSS scores in the pre- and post-experimental phases were 24.75 ( = 7.93) and 22.33 ( = 8.66), respectively (Figure 6). Our Wilcoxon signed-rank test revealed a significant difference the pre- and post-experimental phases ( = 1, = 0.448, ¡ .05, = 0.0746). Experiencing mobile expressive writing through the system for three weeks thus significantly reduced the stress.
@@
The average scores of immediate stress change were -0.40 ( = 0.46), -0.49 ( = 0.55), and -0.57 ( = 0.48) under the None, Passive, and Interactive conditions, respectively (Figure 6). Our Friedman test did not reveal any significant difference ( = 2.64, = .27, = 0.25).
@@
We also examined the ratings of post-experiment cathartic effect. Figure 6 shows the average ratings of all the three conditions: 3.44 ( = 0.97), 3.75 ( = 1.11), and 3.5 ( = 1.13) under the None, Passive, and Interactive conditions respectively. Our Friedman test did not reveal a significant difference ( = 2.14, = .34, = 0.21).
@@ 6.2. Qualitative Results for Stress-relieving Effect
Our participants mentioned the cathartic effects as the advantages of Passive and Interactive conditions though our analysis found differences between them. The major difference emerged from the difference in agency of elimination. In the Interactive condition, the participants gained the cathartic effect as they exploded the letters by themselves. Seven participants felt like they were relieving their stress on their own: “I myself had exploded the stress, and it felt refreshing.” (PB32). The presence of interactivity was influencing agency in stress relief, and our participants pointed out the lack of agency as the disadvantage of the Passive condition. They felt they were “just like observing” (PB28) and not having control. In the visual feedback of stress disappearance, giving control over the vanishing provided participants with a sense of active control over the process, and it enhanced the cathartic effect.
Our participants saw the cathartic effects of the Passive condition by gaining an opportunity to take a distance from the experienced stress. Five participants mentioned laughter as the advantage of the Passive condition: “It was good that it was simply amusing and made me laugh. I thought it was especially good when I was feeling stressed out and bothered.” (PB11) This explicit visualization of the explosion of the externalized stress-related emotion also made the participants feel refreshed or calm down: “Seeing the stress words explode and disappear made me feel a bit amusing in a ‘deserve it!’ kind of way, and I appreciated that it made me feel calm.” (PB30).
@@
However, our participants saw the value of this condition outside the cathartic effect. Participants commented that the system was actively listening to their expressions in this condition. They appreciated the immediate replies from our system, which may relate to this effect. The reply message333The message used in the second user study was “Thank you for sharing this stressful event. We understand that some of this information may be difficult to share, but we greatly appreciate your cooperation.” was pre-defined and used across all the conditions. Some of our participants considered this reply as feedback to them as it offered them a sense of confirmation that they were being heard: “It was beneficial to be able to directly describe the accumulated stress of the day to the chatbot. The immediate response from the chatbot made me feel like it was actively listening to my thoughts.” (PB26)
Organizing Thoughts
Another benefit perceived by our participant was to help participants organize their thoughts. Seven participants mentioned that they could organize their thoughts and objectively recognize and contemplate their stress: “I was able to reflect on the events that caused stress and think about them logically. It was beneficial to calmly reconsider the origins of feelings of aversion and anger towards others. It was beneficial to calmly reconsider the origins of feelings of aversion and anger towards others.” (PB11) This result suggests that the absence of the visual feedback may encourage reflection of participants’ negative feeling, promoting another benefit of expressive writing.
@@ 6.4. Qualitative Results for Preferences
6.4.1. Scenario-dependent Use of Feedback Designs
Our participants shared their views on different use scenarios for each design. They commented that the None condition could be useful when they simply wanted someone else to hear their complaints. Another notable use scenario was for reflecting on their feelings through writing or leaving records for future reflections: “I want to use it (the None condition) when I can’t talk to others about something, but it’s hard for me to keep it to myself, and I want to be heard, even if it’s by a robot.” (PB29)
With respect to the Passive condition, our participants expressed that they would use it when the stress originated from others. Other participants also explicitly mentioned that they would rather avoid using the Passive design when the cause of their stress was associated with themselves: “I want to use it (the Passive condition) when I feel stressed about others. I would not use it for stress within myself because I don’t feel it is very effective.” (PB13)
@@ 6.4.1. Scenario-dependent Use of Feedback Designs
In summary, our qualitative results highlight different possible use scenarios of the three visual conditions that the quantitative results did not uncover. Our participants considered that the None condition would be rather useful for seeking mental cool-down. Their comments also clarified the similarities and differences between the Passive and Interactive conditions. Both feedback designs were considered better suited for stress stemming from others, while the Interactive design can be effective for supporting the relief of intense stress.
6.4.2. Further Explorations of Visual Feedback Designs
Our participants offered various ideas on other possible visual feedback designs. They are categorized into three design philosophies: designs that offer immediate destruction; designs that offer gradual disappearance; and designs that offer detachment. Designs to show immediate destruction included braking blocks and letting virtual characters eat.
Designs to show gradual disappearance are visual effects where expressed negative experience would slowly disappear, such as a fade-away effect of the written text. Our participants commented that such effect may offer more calm feeling for stress relief: “I felt a sense of calm with a video that had a fading text effect, and I would like to continue using it. I thought the explosion effect was too exciting for stress relief and not very effective.” (PB30)
Designs to show detachment intend to allow users to simply leave their negative experience outside themselves. This design philosophy matches with the None condition, which may promote reflection. Our participants suggested a calm visual feedback that is not directly related to the written experience, such as flowers blooming up and animals playing around: “I think something soothing or relaxing, like animals, would be good.” (PB8)
@@ 7. Discussion
7.1. RQ1: Advantages of Keyboard-based Text Entry for Stress-related Emotional Expression
Consistent with previous research in psychological literature (Murray and Segal, 1994), our results from the first study confirmed that both speech and text input produce similar immediate cathartic effects. This result helps to understand the advantage of therapy interviews to expressive writing (Donnelly and Murray, 1991; Murray et al., 1989), which may lie in not the mode of expression (speech vs. written) but the presence of the listener.
Our results also showed that Keyboard-based text entry had an advantage in usability and preference when compared to Voice message input. This contradicts the previous studies (Kim et al., 2018; Luo, 2022; Rzepka et al., 2022; Kocielnik et al., 2018) indicating the superiority of speech input to keyboard input in usability and preference in other use contexts. Our qualitative results suggest two possible explanations: the confidentiality and the affordances of the keyboard-based entry are a better fit for expressive writing that can involve private matters.
Stress-related expression can contain highly-private content. In recording voice messages, our participants spoke their thoughts aloud in various settings, including in public spaces. This privacy threat has been a major concern of voice assistants and depends on the type of information and the location (Easwara Moorthy and Vu, 2015). However, these concerns may not be particularly serious in case where users utter non-private content (e.g., commands (Kim et al., 2018) or short journaling (Luo, 2022)) in private situations. As discussed in Section 4.1, our participant recognized the value of keyboard-based input even in private settings as it allowed participants to relax, write at their own pace, and avoid additional strain or stress. Our findings thus suggest that effects by input modality differences can form important application-dependent effect, stimulating further application-focused research
Our results also suggested that keyboard-based text entry may have helped the cognitive restructuring with its affordances. The keyboard-based condition was preferred because of its persistence and editability. These two affordances (Treem and Leonardi, 2013) may have enhanced the effective process of expressive writing: exposure and cognitive processing. Typing can make the explanation of one’s stress visible and reviewable persistently to oneself (Clark and Brennan, 1991). This may have worked as exposure to their past stressful events, a significant process in expressive writing (Sloan and Marx, 2004). Editability offers a capability to recraft and establish an ideal explanation (Walther, 1993). Editing the text during expressive writing may encourage people to reconsider our thoughts and reconstruct perceptions, known as cognitive processing, another important process to change our perspectives positively (Pennebaker and Chung, 2007). Keyboard-based text entry was thus able to successfully offer desirable features for mobile expressive writing.
7.2. RQ2: Visual Design Preferences Affected by Perceived Cathartic Effects, Sense of Agency, Required Effort, and Sense of Self-Denial
Our results in the second study unveiled different benefits and possible use scenarios of the three visual feedback conditions. This result is not fully aligned with prior work showing the advantages of performing post-writing tearing that create a feeling of destruction (Briñol et al., 2013; Grieger et al., 2021). One possible explanation for this contradiction is that stress relates to diverse conditions and emotions. According to our results, the cause of stress, particularly whether it comes from users themselves or others, was one of the determining factors to the perceived appropriateness of feedback designs. Previous studies only examined emotion about specific topics, such as one’s body or Mediterranean diet (Briñol et al., 2013) or envious experience (Soesilo et al., 2021). Previous literature also reported the influence by the source of emotion. Briñol et al. found that negative emotions related to participants’ bodies persisted even after tearing up their written thoughts, while negative emotions about the Mediterranean diet disappeared (Briñol et al., 2013). Grief therapy research has examined the relationship between the type of emotion and the way of disposal (Sas et al., 2016). Dynamic disposal (e.g., burning) is suitable for anger, but open or covert disposal (e.g., drifting or burying) is rather appropriate for sadness or regret. Our results suggest that future systems should combine different feedback designs to accommodate people’s various types of stress and emotions, offering an important direction for future research.
Participants also found that the None design was still useful for calmly reflecting on their emotions. This might be another explanation for our quantitative results not showing differences in the perceived cathartic effect. The reply to the expressive writing may have functioned as active listening. Active listening is one of the key skills in counseling (Hutchby, 2005) and helps speakers be more emotionally mature and open to their experiences (Rogers and Farson, 1957). Our results suggest that future research on stress management systems should consider purposeful absence of visual feedback to encourage users’ focus on their narrative and promote cognitive restructuring.

### v3

@@ 5.1.2. Dependent Variables for Stress-relieving Effect
Similarly to our first study, we collected immediate stress change and post-experiment cathartic effect. Additionally, we gathered the following quantitative data to assess the effects on mental health conditions beyond subjective impressions.
@@
We used the Perceived Stress Scale (PSS) (Cohen et al., 1994) to measure changes in participants’ stress levels over time. This tool has 14 questions on a five-point Likert scale (0: never to 4: very often), assessing the past week’s experiences. Weekly scores were calculated by subtracting the initial score from the later score related to the condition experienced that week.
@@ 5.1.3. Dependent Variables for Usability and Preference
We collected ease of use, mental workload, and preference using the same questions as Study 1. We also collected effort, by asking participants to rate the statement, “This feedback requires effort,” on a five-point Likert scale (1: “Not at all” – 5: “Strongly agree”), based on prior research comparing usability of digital coping methods (Howe et al., 2022).
@@ 5.2. Procedure
The process in the onboarding and post-experimental sessions was the same as in Study 1. In this study, we explained our participants the objective of our study to examine the effects of three feedback designs at this time. The pre-experiment questionnaire included the questions from PSS for long-term stress change. The onboarding required approximately 30 minutes to complete on the first day. We created six groups of participants with considering the balance of gender, age, and occupations to fully counter-balance the order of the presentations of the three feedback conditions.
Participants are under each of the visual feedback conditions for each week of the experiment (Day 2–8, Day 9–15, and Day 16–22). On the mornings of days 9 and 16, we asked participants to fill out PSS through Google Forms to track their long-term stress level trends.
We conducted post-experiment questionnaires (Table 4 in Appendix B) on the final day, taking approximately 30 minutes. They included PSS questions and items on post-experiment cathartic effect, effort, ease of use, mental workload of each feedback condition, and the preference of feedback conditions. We gathered qualitative data through open-ended questions on their views about the pros and cons of each feedback condition, reasons for their preferences, preferred usage times, and desired feedback designs, aiming to better understand the quantitative results.
We offered all participants approximately 21 USD in the local currency for the completion of the study. In consideration of the longer study duration and increased expected workload compared to Study 1, we provided a higher level of compensation.. We performed data analysis in the same way as we did in our first study.
The ethical considerations for Study 2 followed the same procedures described in Study 1. The only difference was that Study 2 did not include deception or a subsequent debriefing.
@@ 5.3. Participants
We used the same crowdsourcing service as Study 1. We also followed the same procedure to choose participants. As a result, we recruited 36 participants (18 women, 18 men; PB1 – 36) in their 20s to 60s (Table 2). All participants self-claimed that they did not know of the existence of our first study. Additionally, we verified that none of the participants had usernames matching those from Study 1.
@@ 6. Study 2 Results
In this section, we present the results of the second study, covering stress reduction and usability through both quantitative and qualitative analyses. We collected 550 expressive writing entries with immediate stress change data, and the response rate was 99.4%. Among the 550 expressive writing entries, 186, 183, and 181 were under the None, Passive, and Interactive conditions, respectively. The total length of responses across 16 open-ended questions averaged approximately 557 characters per participant, ranging from 181 to 1882 characters, with a response rate of 100%.
@@ 6.1. Quantitative Results for Stress-relieving Effect
The average difference between pre- and post-experimental PSS scores were 0.39 ( = 5.23), -0.58 ( = 5.03), and -2.22 ( = 7.06) in the None, Passive, Interactive conditions, respectively (Figure 6). Our Friedman test did not reveal any significant difference ( = 1.22, = .54, = .117). For the Interactive condition, our Wilcoxon signed-rank test found a significant difference between pre- and post-experimental phases ( = 1, = 1.05, ¡ .05, = .175). For conditions None and Passive, the Wilcoxon signed-rank test showed no significant differences pre- and post-test (None: = 1, = 0.267, = .49, = 0.0445; Passive: = 1, = 0.0393, = .25, = 0.00655).
However, using our system for three weeks resulted in a significant stress reduction. The average PSS scores in the pre- and post-experimental phases were 24.75 ( = 7.93) and 22.33 ( = 8.66), respectively (Figure 6). Our Wilcoxon signed-rank test revealed a significant difference the pre- and post-experimental phases ( = 1, = 0.448, ¡ .05, = 0.0746). Experiencing mobile expressive writing through the system for three weeks thus significantly reduced the stress.
@@
The average scores of immediate stress change were -0.40 ( = 0.46), -0.49 ( = 0.55), and -0.57 ( = 0.48) under None, Passive, and Interactive conditions, respectively (Figure 6). We calculated immediate stress change from participants’ average ratings on each feedback. The Friedman test showed no significant difference ( = 2.64, = .27, = 0.25).
@@
Figure 6 shows the average ratings of all the three conditions: 3.44 ( = 0.97), 3.75 ( = 1.11), and 3.5 ( = 1.13) under the None, Passive, and Interactive conditions respectively. Our Friedman test did not reveal a significant difference ( = 2.14, = .34, = 0.21).
@@ 6.2. Qualitative Results for Stress-relieving Effect
From our thematic analysis, we revealed that each condition provided the cathartic effect differently. The variety in the existence of visual feedback and agency created different outcomes.
@@
Participants felt the cathartic effects in Passive and Interactive conditions, yet differences arose in agency of elimination. In Interactive condition, exploding letters themselves provided relief for seven participants: “I myself had exploded the stress, and it felt refreshing.” (PB32). Interactivity influenced stress relief agency, with lack of it being a downside in Passive condition—they felt “just like observing” (PB28) and out of control. Control over visual stress disappearance enhanced the cathartic effect.
Participants experienced cathartic benefits from the Passive condition by distancing from stress. Laughter, noted by five participants, was an advantage of the Passive condition: “It was good that it was simply amusing and made me laugh. I thought it was especially good when I was feeling stressed out and bothered.” (PB11). The clear visualization of stress-related emotion release also refreshed or calmed them: “Seeing the stress words explode and disappear made me feel a bit amusing in a ‘deserve it!’ kind of way, and I appreciated that it made me feel calm.” (PB30).
@@
However, our participants saw the value of this condition outside the cathartic effect. Participants noted that the system actively listened to their expressions and appreciated the immediate responses. The pre-defined reply message333The message used in the second user study was “Thank you for sharing this stressful event. We understand that some of this information may be difficult to share, but we greatly appreciate your cooperation.”, used in all conditions, provided them feedback and confirmation of being heard. “It was beneficial to be able to directly describe the accumulated stress of the day to the chatbot. The immediate response from the chatbot made me feel like it was actively listening to my thoughts.” (PB26)
Participants also found benefit in organizing their thoughts. Seven noted they could recognize and contemplate stress: “I was able to reflect on the events that caused stress and think about them logically. It was beneficial to calmly reconsider the origins of feelings of aversion and anger towards others.” (PB11). This implies that the absence of visual feedback may encourage reflection, enhancing expressive writing’s benefits in a different manner from the conditions with visual feedback.
@@ 6.3. Quantitative Results for Usability & Preference
With the qualitative analysis of usability and preference metrics, we found that the Interactive condition required more effort and mental workload than the Passive condition did. However, the provision of visual feedback did not compromise the usability of the system, and the participants exhibited a neutral preference for the conditions.
@@ 6.4. Qualitative Results for Preferences
6.4.1. Scenario-dependent Preferences
Our qualitative results uncovered different rationales for preferences of each condition. The None condition was found useful when they simply wanted someone else to hear their complaints. Our participants also commented that it could help them reflect on their feelings through writing or leaving records for future reflections: “I want to use it (the None condition) when I can’t talk to others about something, but it’s hard for me to keep it to myself, and I want to be heard, even if it’s by a robot.” (PB29)
Our participants expressed that they would use the Passive condition when the stress originated from others. Other participants also explicitly mentioned that they would rather avoid using the Passive design when the cause of their stress was associated with themselves: “I want to use it (the Passive condition) when I feel stressed about others. I would not use it for stress within myself because I don’t feel it is very effective.” (PB13)
@@ 6.4.1. Scenario-dependent Preferences
Our qualitative results reveal scenarios for the three visual conditions not shown in the quantitative data. Participants found the None condition useful for mental cool-down and relaxation. Comments highlighted that both the Passive and Interactive conditions served better for stress from others. The Interactive design helped to relieve intense stress though subjective workload was higher in the other designs.
6.4.2. Other Potential Feedback Designs
Our participants offered various ideas on other possible visual feedback designs. They are categorized into three design principles: immediate destruction; gradual disappearance; and detachment. While the Passive and Interactive conditions represent immediate destruction, our participants suggested other designs, such as braking blocks and letting virtual characters eat text. Gradual disappearance is visual effects where expressed negative experience would slowly disappear, such as a fade-away effect of the written text. Our participants commented that such effect may offer more calm feeling for stress relief: “A video where the text gradually fades out feels calming, making me want to use it. I thought the explosion effect was too exciting for stress relief and not very effective.” (PB30) Detachment is intended to allow users to simply leave their negative experience outside themselves. This design principle matches with the None condition, which may promote reflection. Our participants suggested a calm visual feedback that is not directly related to the written experience, such as flowers blooming up and animals playing around: “I think something soothing or relaxing, like animals, would be good.” (PB8)
6.5. Summary
In our second study, participants reported a significant reduction in stress levels after completing the experimental sessions. However, we found no significant differences in cathartic effects among the visual feedback conditions. Qualitative analysis further revealed distinct ways each visual feedback influenced users’ perceptions and emotional experiences. In the next section, we discuss these findings in relation to our original research questions and to relevant literature.
@@ 7. Discussion
7.1. Advantages of Keyboard-based Text Entry in Mobile Expressive Writing
We conducted Study 1 to answer RQ1, which highlighted the unique benefits of keyboard-based text entry over voice input. Aligned with previous psychology research (Murray and Segal, 1994), our results showed that keyboard-based and voice inputs provided similar cathartic effects. However, keyboard text entry slightly outperformed voice input in usability and preference. This challenges past findings of speech input’s superiority in different contexts (Kim et al., 2018; Luo, 2022; Rzepka et al., 2022; Kocielnik et al., 2018). Qualitative insights suggest that the affordances of keyboard entry helped ensure confidentiality (Easwara Moorthy and Vu, 2015) but also better supported the cognitive restructuring process (Pennebaker and Chung, 2007) central to expressive writing.
Participants favored keyboard-based text entry for its persistence and editability (Treem and Leonardi, 2013), enhancing two key psychological mechanisms of expressive writing: exposure (Sloan and Marx, 2004) and cognitive processing (Pennebaker and Chung, 2007). Typing enables the externalization of one’s stress in a persistent, reviewable format (Clark and Brennan, 1991), functioning as an exposure to past stressors — an important process in expressive writing (Sloan and Marx, 2004). Moreover, the ability to revise enabled participants to reconstruct their narratives (Pennebaker and Chung, 2007). From the perspective of cognitive processing theory, expressive writing involves a dynamic process of planning, translating, and reviewing, rather than a simple linear flow of thoughts (Flower and Hayes, 1981). In this regard, typed input enabled users to pause, reflect, and revise their expressions (Daiute, 1986), thereby supporting narrative reorganization. In contrast, voice input encouraged a more linear, less editable form of narration (Schalkwyk et al., 2010; Lai and Yankelovich, 2006), which made it harder for participants to revise or redirect their thoughts once expressed. Previous research has similarly shown that typed input on mobile devices offers superior editing capabilities compared to speech input (Cherubini et al., 2009; Luo et al., 2022), which may facilitate deeper cognitive processing through iterative reflection and reinterpretation (Pennebaker and Chung, 2007).
On the other hand, Murray et al. (Murray et al., 1989) and Esterling et al. (Esterling et al., 1994) reported that expressing with voice into a tape recorder included more content reflecting cognitive changes and negative emotions. While using the same modality to express stressful experiences, our study did not show similar outcomes. This comparison proposes that chat interface may bring audience impairment effects (Lord et al., 1987). Some of our participants felt nervous when using voice input, possibly because they were conscious of the presence of the researcher as the recipient. The audience effect is known to influence behavior even when the audience is merely imagined (Higgs and Joseph, 1971), and its impact tends to be stronger when the imagined audience is someone with whom the individual has low involvement (Axsom et al., 1987). Taken together, these findings suggest that in a chat interface, sending a voice message may evoke a strong sense of being monitored while recording. This perceived observation could create pressure that disrupts cognitive processing.
Building on these findings, designing the voice-based mobile interface for expressive writing requires addressing both privacy concerns and the audience impairment effect. To mitigate this, systems should foster a sense of psychological safety. One approach is to reduce an audience’s perceived presence, for example, by making it clear that recordings are stored only locally and not automatically transmitted or by explicitly separating the recording and sending processes to give users more control. Alternatively, designers could embrace the presence of an audience by providing responsive voice feedback, such as affirmations or empathetic cues, similar to how text-based chatbots build affirmative relationships through empathic responses (Fitzpatrick et al., 2017). These strategies highlight the importance of carefully shaping the social dynamics users perceive during voice-based emotional expression.
7.2. Feedback Design Preferences Affected by Multiple Factors
The results of Study 2 to answer RQ2 revealed distinct characteristics and benefits for all three feedback conditions, along with their possible use scenarios. Previous work has shown how post-disclosure interaction helps alleviate negative feelings toward certain topics, e.g., by disposing of the paper on which envious feelings were disclosed, compared to a baseline of doing nothing (Briñol et al., 2013; Grieger et al., 2021; Soesilo et al., 2021). We did not, however, observe a similar effect in our study as there were no significant differences between the conditions concerning stress release. Our results suggest that digital tools like what we developed in this study, while effective, may offer different effects from what is observed between the physical act of discarding a paper into a bin and taking no action, encouraging further research.
By allowing participants to express stress in any context in their lives, our study covered a diverse set of conditions and sources of stress. In our case, particularly whether the stress was self-caused or had something to do with other people, i.e., the source, was one of the determining factors to the perceived appropriateness of feedback designs. Related to this effect of the source, Briñol et al. found that negative emotions related to participants themselves persisted even after tearing up their written thoughts, while negative emotions about others disappeared (Briñol et al., 2013). Our findings help confirm the effect of the source with a broader set of stressful moments people may face in their daily lives, expanding the existing understanding of the effect of expressive writing.
Our study also found that the Passive condition, where the feedback did not provide a sense of agency, induced laughter. The incongruity between expectation and outcomes can generate laughter (Rothbart, 1973). The unexpected, exaggerated explosion provoked laughter, and the laughter overcame the negative emotion (Keltner and Bonanno, 1997). The combination of both objectified emotion and the act of laughter may augment the cathartic effect.
The Interactive condition, where users tapped to trigger the visual disappearance, was perceived as more effective in relieving stress. Participants attributed this effect to their ability to actively control the experience. This sense of agency (Moore, 2016; Alhasani et al., 2022) appeared to amplify the cathartic release of inhibited emotions beyond what the writing process alone could offer (Pennebaker and Beall, 1986). Providing users with opportunities to act at their own pace can enhance their sense of agency and personal involvement, which are known to support emotional regulation. However, our findings also indicate that increased interactivity may introduce additional cognitive or physical workload. To address this trade-off, it may be beneficial to design interventions that offer adjustable levels of agency. For instance, system could allow users to choose between passive and interactive modes, or dynamically adapt the level of interactivity based on contextual factors or intensity of stress. Such flexibility may improve both user engagement and the overall effectiveness of the system.
Participants also expressed that the None design was also valuable for reflecting on their emotions. One possible explanation for this effect is that this condition may have offered an environment of active listening. Active listening, which is one of the key skills in counseling (Hutchby, 2005), supports the speakers to be more emotionally mature and open to their experiences (Rogers and Farson, 1957). Our system sent participants notifications after they submitted their disclosure to confirm its reception, but it may have served as a way to demonstrate listening.
Based on our results, stress-coping interventions should consider the characteristics of stress to improve their acceptability. A growing body of research has developed stress-coping support systems that deliver just-in-time adaptive interventions (JITAI) (Nahum-Shani et al., 2018; Han et al., 2022). Recent adaptive systems often utilize biometric data (e.g., heart rate collected from smartwatches (Han et al., 2022; Wakschlag et al., 2021)) or self-reported data (e.g., stress levels collected via ecological momentary assessment, or EMA (Juarascio et al., 2021; Presseller et al., 2022)) as tailoring variables to determine the optimal timing of interventions. By scaling the level of stress based on these data, these systems have successfully delivered interventions when users are most vulnerable, supporting more informed decision-making. However, an equally important factor—receptivity—has largely been overlooked (van Genugten et al., 2025). In light of our findings, we suggest that effective intervention design should consider not only the intensity of stress but also the orientation of stress when determining the nature and timing of support. In this context, expressive writing, as an activity that involves articulating personal experiences and emotions, may play a dual role: not only as a therapeutic tool, but also as a means of sensing. It can offer valuable insights into the sources and directions of stress, making it a promising component of mobile systems for everyday stress management and adaptive interventions.